Determine the limit when x goes to infinity sqrt(x^(2)+x)-x if it exists.


1
Expert's answer
2021-02-01T07:32:23-0500

"The, initial, form, for, the, limit, is, indeterminate,
\infty-\infty
So, use, the, conjugate.
\begin{array}{l}
\left(\sqrt{x^{2}+x}-x\right)=\frac{\sqrt{x^{2}+x}-x}{1} \\
\frac{\sqrt{x^{2}+x}+x}{\sqrt{x^{2}+x}+x} \\
=\frac{x^{2}+x-x^{2}}{\sqrt{x^{2}+x}+x} \\
=\frac{x}{\sqrt{x^{2}+x}+x}
\end{array}
\lim _{x \rightarrow \infty} \frac{x}{\sqrt{x^{2}+x}+x} has, indeterminate, form,
\frac{\infty}{\infty}, but, we, can, factor, and, reduce.
We, know, that, \sqrt{x^{2}}=|x|, so, for, positive, x
(which, is, all, we, are, concerned, about, for, a, limit, as, x, increases, without, bound,) we, have,
\begin{array}{l}
\frac{x}{\sqrt{x^{2}+x}+x}=\frac{x}{\sqrt{x^{2}} \sqrt{1+\frac{1}{x}+x}} \\
x \neq 0)
\end{array}
(for, all""=\frac{x}{x \sqrt{1+\frac{1}{x}}+x}
x>0)
=\frac{x}{x\left(\sqrt{\left.1+\frac{1}{x}+1\right)}\right.}
=\frac{1}{\sqrt{1+\frac{1}{x}}+1}
\lim _{x \rightarrow \infty} \frac{1}{\sqrt{1+\frac{1}{x}}+1}=\frac{1}{\sqrt{1}+1}=\frac{1}{2}"